OpenStudy (anonymous):

Simplify the equation 4m^4n^4 / 14m

OpenStudy (anonymous):

just follow the rule: \[\frac{a^x}{a^y}=a^{x-y}\]

OpenStudy (anonymous):

\[\frac{4m^4n^4}{14m}=\frac{2}{7}m^{4-1}n^4=~?\]

OpenStudy (anonymous):

\[\frac{ 2m^3n^4 }{ 7 }\]?

OpenStudy (anonymous):

@orion1213

OpenStudy (anonymous):

yup it's correct....